Prepare the cash flow statement from investing activities of Alpha Creative Ltd for the year ended March31, 2019
Particulars Amount
Plant acquired 160000
Claim received for loss of plant in fire 45500
Unsecured loans given to subsidiaries 595000
Interest on loan received from subsidiary companies 72500
•Also give reasons for the classification of above activities as inflow /outflow
1
Expert's answer
2019-12-02T09:18:02-0500

Cash Flow from Investing Activities is the section of a company’s cash flow statement that displays how much money has been used in (or generated from) making investments during a specific time period. Investing activities include purchases of long-term assets (such as property, plant, and equipment), acquisitions of other businesses, and investments in marketable securities (stocks and bonds).


Cash receipts: 45500 + 72500 = 118000.

Cash paid: 160000 + 595000 = (755000).

Net cash flow: 118000 - 755000 = (637000).